### v7

@@ 3 Main results
In the remainder of this paper, the symbol $c$ denotes a generic positive constant that depends only on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{n}$, the constant $C_{F}$ in Eqs. 2.1, 2.2, 2.4 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ given in equations Eq. 1.1 and Eq. 2.11, the regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with the spatial mesh size $h$. The value of $c$ may differ at each occurrence. For any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\|v\|_{\dot{H}^{\alpha,q}}:=\|(-\Delta)^{\alpha/2}v\|_{L^{q}}\quad\text{for all $v\in\dot{H}^{\alpha,q}$}.$
The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$. It is well-established that $\Delta$ has a unique extension as a bounded linear operator from $\dot{H}^{\alpha,q}$ to $\dot{H}^{\alpha-2,q}$ for all $\alpha\in[0,2)$. Specifically, when $\alpha=2$, the space $\dot{H}^{2,q}$ is identical to the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$ and is endowed with a norm equivalent to that of $W^{2,q}(\mathcal{O})$ (see [11, Theorem 9.15]). Moreover, for $\alpha=1$, the space $\dot{H}^{1,q}$ coincides with the standard Sobolev space $W_{0}^{1,q}(\mathcal{O})$, and the two norms are equivalent (see [38, Theorem 16.14]).
The main results of this paper are summarized in the following three theorems.
@@
Assume that $v\in L^{\infty}$. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
@@
Furthermore, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
@@
Assume that $v\in L^{\infty}$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq. 2.11. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{\alpha/p}+h^{2/p}t^{-1/p}\right).$ (3.6)
@@
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution of Eq. 2.11. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
@@
Liu [18] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields
@@
provided that $v\in\dot{H}^{1,2}$. Furthermore, Liu and Qiao [21] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem. Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
@@
  2. 2.
    In Section 4.3, we show that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, assuming $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.53). It can be rigorously demonstrated that this result also holds when $v\in W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This finding is critical for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, it remains an open question whether this conclusion applies when $v\in L^{\infty}$.
@@ 4.1 Preliminary Lemmas
It is a well-established result that for any $q\in(1,\infty)$ and $s\in(0,2)$, the space $\dot{H}^{s,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{s/2}$, equipped with equivalent norms. This result is documented, for example, in Chapter 16 of [38]. The combination of Proposition 1.4 from [22] and Theorem C.4.1 in [12] yields the following continuous embeddings.
@@
Let $0\leqslant\beta<\gamma\leqslant 1$ and $q\in(1,\infty)$. Then the following continuous embeddings hold:
@@
    The space $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
@@
    The space $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
@@
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. For all $t\in(0,T]$, the following estimate holds:
$\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
For the deterministic convolution operator $S_{0}$, defined by Eq. 2.7, we have the following celebrated deterministic maximal $L^{p}$-regularity estimate; see Theorem 6.1.11 in [30] and Corollary 1.14 from [22].
@@
Let $q\in(1,\infty)$. The following inequalities are valid:
@@
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
@@
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q},$
where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [13]), and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
@@
Analogously, employing the growth condition Eq. 2.2 yields
@@
and utilizing the Lipschitz continuous condition Eq. 2.3 results in
@@
For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
$y_{n}(0)=v_{n}$ (4.6b)
@@
$\lVert{y_{n}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{v_{n}}\rVert_{L^{q}}).$ (4.7)
@@
Step 1. Define $G_{n}(t):=e^{t\Delta}v_{n}$ for $t\in[0,T]$. Given that $v_{n}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
$G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
In view of Eq. 4.2 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [18, Theorem 1.1] (see also, [21, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y_{n}$ and
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which in turn implies that $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
Since (4.2) and (4.9) ensure that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
@@
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
@@
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
@@
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact
$y_{n}(t)=v_{n}+\int_{0}^{t}(\Delta y_{n}+y_{n}-y_{n}^{3})(s)\,\mathrm{d}s+\int_{0}^{t}F(y_{n}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{[}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{]}$
@@
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s\Biggr{]}$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\bigg{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
@@
$\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)(\Delta y_{n}-y_{n}^{3})(s)\,\mathrm{d}\mu$
@@
$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 along with the growth condition Eq. 2.2, we derive the following bound:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$

### v8

@@ 3 Main results
In the remainder of this paper, we denote by $c$ a generic positive constant that is independent of the spatial mesh size $h$. It may, however, depend on the final time $T$, the domain $\mathcal{O}$, the constant $C_{F}$ from Eqs. 2.2 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial data $y_{0}$, the shape-regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with $h$, unless otherwise specified. The value of $c$ may differ at each occurrence.
@@
Assume that $y_{0}\in L^{\infty}$ is deterministic. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
@@
Furthermore, for any $p,q\in(2,\infty)$ and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
@@
Assume that $y_{0}\in L^{\infty}$ is deterministic. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq. 2.9. Then, for any $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\right),\quad\forall\epsilon>0.$ (3.6)
@@
Assume that $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ is deterministic. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution of Eq. 2.9. Then the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2-\epsilon},\quad\forall p\in(2,\infty),\,\forall\epsilon>0.$ (3.8)
@@
Liu [20] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to the model problem Eq. 1.1 yields
@@
provided that $y_{0}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};\dot{H}^{1,2})$, where $p\in[2,\infty)$. Furthermore, Liu and Qiao [23] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem. Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
@@ 4.1 Preliminary Lemmas
It is a well-established result that for any $\alpha\in(0,2)$ and $q\in(1,\infty)$, the space $\dot{H}^{\alpha,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{\alpha/2}$, equipped with equivalent norms. This result is documented, for example, in Chapter 16 of [37]. The combination of Propositions 1.3 and 1.4 and Theorem 4.17 from [24] and Theorem C.4.1 in [15] yields the following continuous embeddings.
@@
Let $0\leqslant\alpha<\beta\leqslant 2$ and $p,q\in(1,\infty)$. Then the following continuous embeddings hold:
@@
    If $\gamma>(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then $\dot{H}^{\gamma,q}$ is continuously embedded in the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$.
@@
    If $\gamma<(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$ is continuously embedded in $\dot{H}^{\gamma,q}$.
@@
Let $\alpha\in[0,\infty)$ and $q\in(1,\infty)$. For all $t>0$, the following estimate holds:
$\|(-\Delta)^{\alpha}e^{t\Delta}\|_{\mathcal{L}(L^{q})}+t\Big{\|}(-\Delta)^{\alpha}\frac{\mathrm{d}}{\mathrm{d}t}e^{t\Delta}\Big{\|}_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$
For the convolution operator $S_{0}$, defined by Eq. 2.5, we have the following celebrated maximal $L^{p}$-regularity estimate; see Theorem 4.2 in [36] and Corollary 1.14 from [24].
@@
Let $q\in[2,\infty)$. The following inequalities are valid:
@@
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}f_{n}(v),\quad\forall v\in L^{q}.$
@@
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}x\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}1+\lvert{v}\rvert^{2}+\lvert{\nabla v}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q},$
where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [16]), and step $(\mathrm{ii})$ relies on conditions Eqs. 2.3 and 2.2. Using the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}x\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
@@
Analogously, employing the condition Eq. 2.2 yields
@@
and utilizing the condition Eq. 2.3 results in
@@
For each $n\in\mathbb{N}$ and $y_{0}^{(n)}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, which is deterministic, the stochastic Allen-Cahn equation
$\mathrm{d}y^{(n)}(t)=\left[\Delta y^{(n)}+y^{(n)}-(y^{(n)})^{3}\right](t)\,\mathrm{d}t+F(y^{(n)}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
$y^{(n)}(0)=y_{0}^{(n)}$ (4.6b)
@@
$\lVert{y^{(n)}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}).$ (4.7)
@@
Step 1. Define $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for $t\in[0,T]$. Given that $y_{0}^{(n)}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
$G^{(n)}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
In view of the estimates for $F$ in Eqs. 4.2 and 4.4 and the fact that $y_{0}^{(n)}\in\dot{H}^{1,2}$ is deterministic, we use [20, Theorem 1.1] (see also, [23, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y^{(n)}$ and
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which in turn implies that $y^{(n)}-(y^{(n)})^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
Since (4.2) and (4.9) ensure that $F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
@@
$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
@@
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
@@
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact, $\mathbb{P}$-almost surely,
$y^{(n)}(t)=y_{0}^{(n)}+\int_{0}^{t}\Delta y^{(n)}(s)+y^{(n)}(s)-(y^{(n)}(s))^{3}\,\mathrm{d}s+\int_{0}^{t}F(y^{(n)}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [8, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{[}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}-\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}\Big{]}$
@@
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lvert{f_{k}(y^{(n)}(s))}\rvert^{2}\,\mathrm{d}x\,\mathrm{d}s\Biggr{]}$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\bigg{[}\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)f_{k}(y^{(n)}(s))\,\mathrm{d}x\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
@@
$\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)[\Delta y^{(n)}(s)-(y^{(n)}(s))^{3}]\,\mathrm{d}x$
@@
$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing condition Eq. 2.2, we derive the following bound:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$